OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. 2x + y = −4 y = 3x + 2 negative 6 over five comma negative 8 over 5 negative 8 over 5 comma negative 6 over 5 negative 5 over 6 comma negative 11 over 5 negative 11 over 5 comma negative 6 over 5

pooja195 (pooja195):

Plug equation 2 into equation 1 (in place of y since they tell u what y equals) 2x+3x+3=-4 Now you can solve for x by yourself. I can guide.

OpenStudy (jchick):

sorry the rest of it didn't go

OpenStudy (anonymous):

@jchick rest of what

OpenStudy (anonymous):

@pooja195 can u help

pooja195 (pooja195):

I did.

OpenStudy (anonymous):

but solving for x still doesn't make my answer very clear to me

OpenStudy (anonymous):

@jchick do u have the answer? :c

OpenStudy (plasmataco):

2x+3x+3=-4

OpenStudy (anonymous):

so a b c or d

OpenStudy (plasmataco):

5x+3=-4

OpenStudy (plasmataco):

5x=-7

OpenStudy (plasmataco):

X=-7/5

OpenStudy (anonymous):

openstudy is a tool to help learn not be given answers. ask questions and get and explanation... but find the answer on your own...

OpenStudy (plasmataco):

Sry, -6/5

OpenStudy (plasmataco):

Now, substitute -6/5 in with x in the first equation

OpenStudy (plasmataco):

If you do the math you get-8/5 so a

OpenStudy (anonymous):

rolo. look at the last answer... If you do the math you get-8/5 so a

OpenStudy (anonymous):

i just closed the question lol

OpenStudy (anonymous):

oh i saw it got bumped to le top